OpenStudy (anonymous):

Mr. Berg, who works twice as fast as Mr. Sloc, receives an hourly rate of pay 1.5 times as much as Mr.Sloc. An efficiency expert calculates that an article priduced by Mr.Berg has a labor cost of 12 cents. What is the labor cost, in cents, of an article, produced by Mr. Sloc?

OpenStudy (anonymous):

produced*

OpenStudy (anonymous):

B=2s for time length x B=1.5s for pay Im assuming that the article was produced in an hour which would be the variable x. B=12 which should be plugged into the equation for pay 12/1.5=Sloc's hourly pay. Since Sloc works for 30 minutes which is B/2, we have to divide pay by 2 per hour or however long the article takes. 12/2=6 cents Any other variable should work

OpenStudy (anonymous):

i got 6 cents too but the answer key says 16.

OpenStudy (anonymous):

oh i see what i did wrong. s=2B not the other way around.

OpenStudy (anonymous):

so sloc would take an hour and berg would take 30 minutes. His hourly pay would be 24 cents. 24/1.5=16

OpenStudy (anonymous):

im terribly sorry for that errror

OpenStudy (anonymous):

Np and thank you for explaining :)